1/2

Thomas Moore,

’T is the Last Rose of Summer

kun ok-liniaj strofoj

’T is the last rose of summer Left blooming alone; All her lovely companions Are faded and gone; No flower of her kindred, No rosebud is nigh, To reflect back her blushes, To give sigh for sigh.

I’ll not leave thee, thou lone one! To pine on the stem; Since the lovely are sleeping, Go sleep thou with them. Thus kindly I scatter Thy leaves o’er the bed, Where thy mates of the garden Lie scentless and dead.

…

2/2

So soon may I follow When friendships decay, And from Love’s shining circle The gems drop away. When true hearts lie wither’d And fond ones are flown, Oh! who would inhabit his bleak world alone?

Verkinto de tiu ĉi Angla poemo estas Thomas Moore (∗1779-05-28 – †1852-02-25).

Arg-531-1058 (2010-03-28 21:25:21)

Mi, Manfred Retzlaff, kopiis tiun poemon el la retejo http: // www. franks. org/ fr01066. htm . Pri la verkinto, la irlanda poeto Thomas Moore, vidu la vikipedian retejon http: // en. wikipedia. org/ wiki/ Thomas_Moore .

http://www.poezio.net/version?poem-id=531&version-id=1058